<<TableOfContents(3)>>

= Agenda/Minutes =
[[https://zoom.us/j/203743987|Zoom Meeting]]

----------
= CRY+VIS Interface Meeting =
Date: 2023/7/14 13:00-13:40

Participants: Takahashi, Miyoki, Yokozawa, Sato, Akutsu, Mitsuhashi, Node, Ushiba, Ikeda, Aso, Hirata

[[https://paper.dropbox.com/folder/show/AcceptanceTest-e.1gg8YzoPEhbTkrhvQwJ2zzo3dHVh5sJGieDgbm2n3h60mGqotBeE|Acceptance check document]]

== Suspensions ==
 * Type-A: Offloaded the BF GAS in IY.
 * Type-B: 
 * Type-Bp: 
 * Type-C: Health check of OMMT1/2/OSTM suspensions.

== Other items ==
 * Oplev: 
 * ACC: Test of the IP control with ACCs in IY.
 * Commissioning: 
 * Vacuum system: Opened the OMC chamber. Leak test in EYC and EXV.
 * Heater control: 
 * TMS: 

== Post O4a ==
 * Measurements of the DAC noise: Every Friday from 7/7.
 * Measurements on the sensor correction in SR2: 1 night.
 * Optimization of the damping control in PRs: Measurements 2 days, Test 3 days.
 * IP control with the ACCs: Postponed the long-term test in IY.
 * In-vacuum accelerometer for IOO: Installed the geophone pod into the OMC chamber on 7/10.
 * Hight adjustments of Type-A in the cooling: Confirm [[https://paper.dropbox.com/doc/Hight-adjustment-in-cooling--B7ii3omB81JIm6khnBt_7fcaAQ-swWQ5h6M6lSs8yLp5D6dA|the procedure]].
 * Check of the length Oplev: 
 * Noise budget of the suspensions: Identified the peak frequencies. Peaks in BS don't match the model.
 * Temperature control for the Oplev: Preparing for the heater controller.

== Discussion ==